OpenStudy (sissyedgar):

A cruise ship can cover 17 nautical miles in 306 minutes. How many nautical miles will it travel in 180 minutes? A. 12 B. 14 C. 8 D. 10

OpenStudy (sissyedgar):

@Holly00d1248

OpenStudy (ilovecake):

C.8 I think

OpenStudy (anonymous):

17/180

OpenStudy (ilovecake):

?

OpenStudy (mathstudent55):

Use a proportion.

OpenStudy (ilovecake):

C.8

OpenStudy (mathstudent55):

17 is to 306 as x is to 180. 17:306 = x:180 \(\dfrac{17}{306} = \dfrac{x}{180} \) Now cross multiply and solve.

OpenStudy (mathstudent55):

Answer is not 8.

OpenStudy (sissyedgar):

UMM. 17*180=3060

OpenStudy (sissyedgar):

X*306=...

OpenStudy (sissyedgar):

IDK

OpenStudy (mathstudent55):